Q1. A firm’s production function is , where L denotes the size of the workforce. Find the value of MPL in the case when: 

(a) L=1, (b) L=10, (c) L=100, (d) L=1000

Does the law of diminishing marginal productivity apply to this particular function?


Q2. Show that the price elasticity of demand is constant for demand functions of the form


where A and n are positive constants.



1
Expert's answer
2021-05-25T17:24:46-0400

Q1.

Given information: 

"Q = 50L - 0.01L^2"

The formula for MPL "= \frac{\delta Q}{\delta L}"

"=50 - 0.02L^{2-1}"

"=50 - 0.02L"


When L = 1 

MPL = 50 - 0.02

    = 49.98units

When L = 10

MPL = 50 - 0.02(10)

    = 50 - 0.2

    = 49.8 units

When L = 100

MPL = 50 - 0.02(100)

    = 48 units 

When L = 1000

MPL = 50 - 0.02(1000)

    = 30 units

"fLL =\frac{\delta MPL}{\delta L} = -0.02"

fLL < 0


When L rises the MPL decreases from 49.98 to 30units.

The law of diminishing marginal productivity is applied in this function.



Q2.

If the form of the demand function is "Q=AP^{-n}" where A and n

 are positive constants, the elasticity of demand is 𝑐

Therefore the elasticity of demand is constant.